OpenStudy (anonymous):

f(x) = x2 - 16 and g(x) = x+4. Find f/g and its domain.

OpenStudy (anonymous):

A. x - 4; all real numbers except x \[\neq-4\] B. x + 4; all real numbers except x \[\neq-4\] C. x + 4; all real numbers except x \[\neq4\] D. x - 4; all real numbers except x\[\neq4\]

OpenStudy (xmoses1):

I am not good at math :\

whitemonsterbunny17 (whitemonsterbunny17):

I'm soooo sorry but I'm not good at math either. :s @suckerofmath can you help? c:

OpenStudy (anonymous):

I have no idea sorry.

OpenStudy (xmoses1):

She is intelligent in an inarticulate manner ^.^ Just kidding Tea :p

OpenStudy (anonymous):

Lol ohh wow

OpenStudy (anonymous):

So first we know that a^2 - b^2 = (a -b) * (a + b) ==> x^2 - 16 = x^2 - 4^2 = (x+4) * (x - 4) Since we are finding domain of f/g = (x+4)*(x - 4)/(x+4) = (x - 4) So Domain of (x - 4) is ??? this function f/g is exist for all real values therefore Domain is R

OpenStudy (anonymous):

I have no idea here

OpenStudy (anonymous):

Okay We are finding domain of f/g = x^2 - 16 /x+4 right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

& we know that a^2 - b^2 = (a -b) * (a + b) So we can write x^2-16 as =x^2 - 4^2 = (x- 4) * (x + 4) okay ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So we have got f/g = (x + 4) * (x - 4) / (x +4) right ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Now we cancel common factor fro numerator & denominator So we will get f/g = (x + 4)*(x-4)/(x+4) = (x - 4) Ok ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Answer is D ?

OpenStudy (anonymous):

Now we have f/g = (x - 4) & we are finding domain Domain is the set of values for which function is defined or function make sence So Here at any value of x our function f/g is defined , that is if we put any real number instead of x we get some answer that is function is defined at all real values Hence Domain = R

OpenStudy (anonymous):

Is D the answer ?

OpenStudy (anonymous):

no please check it the function is undefined when denominator is zero our original function f / g = x^2 - 16 / (x + 4) so at what value denominator become zero ?

OpenStudy (anonymous):

A ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u r welcome :)

OpenStudy (anonymous):

#Shay17 #tgawade was the answer to this problem A? I'm so lost right now